### v1

@@ 4   Conclusion
In summary, we have successfully grown epitaxial Sm2Co17 films by MBE and revealed, for the first time, the the layer- and momentum-dependent electronic structure for a typical REPM. Our results unveil that the Co $3d$ electrons form flat bands near $E_{F}$, one at $\sim$ –300 meV and another right at $E_{F}$, which arise from the orbital-selective destructive interference in a kagome-honeycomb lattice and correlation effects, respectively. Our work further reveals that Sm3+ $4f$ states are far away from $E_{F}$ and exhibit an anomalous temperature evolution, due to the $3d$-$4f$ interaction and competition with strong FM. These findings can be reasonably captured by theoretical calculations, thereby paving the way for a fundamental understanding of REPMs. The spectroscopic insight also hints on alternative approaches to search for high-performance REPMs, through careful design/engineering of crystal structures and associated flat bands. Finally, our results also suggest new routes to explore emergent phenomena associated with flat bands near $E_{F}$ [74, 59, 75], by generating destructive interference in kagome-honeycomb lattices in combination with strong correlation effects.
This work is supported by the National Key R&\&&D Program of China (Grant No. 2022YFA1402200, No. 2023YFA1406303, No. 2022YFA1403202), the National Science Foundation of China (No. 12174331, No. 12350710785), the Key R&\&&D Program of Zhejiang Province, China (2021C01002), the State Key project of Zhejiang Province (No. LZ22A040007) and Zhejiang Provincial Natural Science Foundation of China (Grant No. LGG22E020006). We thank Prof. Frank Steglich, Prof. Julio Larrea, Prof. Zhentao Wang, Prof. Zheng Liu, Prof. Michael Smidman, Prof. Jiaying Jin, Prof. Yanwu Xie, Prof. Zhicheng Zhong, Prof. Xin Lu, Prof. Jiefeng Cao, Dr. Bodry Tegomo Chiogo, Dr. Ruiwen Xie, Dr. Shuaishuai Yin, Dr. Ge Ye and Ms. Xueshan Cao for discussions and help on experiments.

### v2

@@ 5   Conclusion
In summary, we have successfully grown epitaxial Sm2Co17 films by MBE and revealed, for the first time, the the layer- and momentum-dependent electronic structure for a typical REPM. Our results unveil that the Co $3d$ electrons form flat bands near $E_{F}$, one at $\sim$ –300 meV and another right at $E_{F}$, which arise from the orbital-selective destructive interference in a kagome-honeycomb lattice and correlation effects, respectively. Our work further reveals that Sm3+ $4f$ states are far away from $E_{F}$ and exhibit an anomalous temperature evolution, due to the $3d$-$4f$ interaction and competition with strong FM. These findings can be reasonably captured by theoretical calculations, thereby paving the way for a fundamental understanding of REPMs. The spectroscopic insight also hints on alternative approaches to search for high-performance REPMs, through judicious design/engineering of crystal structures and associated flat bands. Finally, our results also suggest new routes to explore emergent phenomena associated with flat bands near $E_{F}$ [76, 60, 77], by generating destructive interference in kagome-honeycomb lattices in combination with strong correlation effects.
This work is supported by the National Key R&\&&D Program of China (Grant No. 2022YFA1402200, No. 2023YFA1406303, No. 2022YFA1403202), the National Science Foundation of China (No. 12174331, No. 12350710785), the Key R&\&&D Program of Zhejiang Province, China (2021C01002), the State Key project of Zhejiang Province (No. LZ22A040007), Zhejiang Provincial Natural Science Foundation of China (Grant No. LGG22E020006), Centro Nacional de Desenvolvimento Científico e Tecnológico (CNPq) Grant No. 404312/2024-1 and and São Paulo Research Foundation (FAPESP) Garnt No. 2018/08845-3. We thank Prof. Frank Steglich, Prof. Zhentao Wang, Prof. Zheng Liu, Prof. Michael Smidman, Prof. Jiaying Jin, Prof. Yanwu Xie, Prof. Zhicheng Zhong, Prof. Xin Lu, Prof. Jiefeng Cao, Dr. Bodry Tegomo Chiogo, Dr. Ruiwen Xie, Dr. Shuaishuai Yin, Dr. Ge Ye and Ms. Xueshan Cao for discussions and help on experiments.
@@
Supporting Information The supporting information is available online at http://phys.scichina.com and https://link.springer.com. The supporting materials are published as submitted, without typesetting or editing. The responsibility for scientific accuracy and content remains entirely with the authors.